# On RoCK blocks of double covers of symmetric and alternating groups and the refined Broué conjecture

### v1

@@ 4 On Rickard equivalences and splendid complexes
4.1.
Let $R$ be a commutative ring and $G$ a finite group. A finitely generated $R$-algebra AA is called $G$-graded (resp. fully $G$-graded) if, as an $R$-module, AA is a direct sum $A=\oplus_{g\in G}A_{g}$ satisfying $A_{g}A_{h}\subseteq A_{gh}$ (resp. $A_{g}A_{h}=A_{gh}$) for all $g,h\in G$; if $A^{\times}\cap A_{g}\neq\varnothing$, then $A$ is called a cross product (of $A_{1}$ and $G$). Let $A$, $B$ and $C$ be fully $G$-graded $R$-algebras. A finitely generated AA-$B$-bimodule $M$ is called $G$-graded if, as an $R$-module, $M$ is a direct sum $M=\oplus_{x\in G}M_{x}$ satisfying $A_{g}M_{x}B_{h}=M_{gxh}$ for all $g,h,x\in G$. Hence with this terminology, if $G=\mathbb{Z}/2\mathbb{Z}$, then $A$ and $B$ are exactly $R$-superalgebras and $M$ is exactly an $(A,B)$-bisupermodule. Following [26], we denote by AA-Gr-$B$ the categroy of $G$-graded finitely generated AA-$B$-bimodules and grade preserving AA-$B$-bimodule homomorphisms. Let $K^{b}(\mbox{$A$-Gr-$B$})$ denote the homotopy category of bound complexes over AA-Gr-$B$; see e.g. [23, page 117] for homotopy categories. Let $M$ be a $G$-graded AA-$B$-bimodule and $N$ a $G$-graded $B$-$C$-bimodule. Then $M\otimes_{B}N$ is a $G$-graded AA-$C$-bimodule with $(M\otimes_{B}N)_{x}=\sum_{gh=x}M_{g}\otimes_{B}N_{h}$. Hence for an object $X$ of $K^{b}(\mbox{$A$-Gr-$B$})$ and an object $Y$ of $K^{b}(\mbox{$B$-Gr-$C$})$, the complex $X\otimes_{B}Y$ is an object of $K^{b}(\mbox{$A$-Gr-$C$})$.
Let $A$ and $B$ be symmetric $R$-algebras. If $X$ is a bounded complex of AA-$B$-bimodules that are finitely generated projective as left $A$-modules and as right $B$-modules, such that $X\otimes_{B}X^{*}$ and $A$ are homotopy equivalent as complexes of AA-AA-bimomdules, and that $X^{*}\otimes_{A}X$ and $B$ are homotopy equivalent as complexes of $B$-$B$-bimodules. Then we say that $A$ and $B$ are Rickard equivalent, and $X$ is called a Rickard complex of AA-$B$-bimodules. Here $A$ (resp. $B$) denotes the complex with only one nonzero term $A$ (resp. $B$) in degree 0, and we adopt [23, Definition 1.17.9] for the definition of tensor product of complexes.
Let $A$, $B$ and $C$ be fully $G$-graded symmetric $R$-algebras. If there is a complex $X\in K^{b}(\mbox{$A$-Gr-$B$})$ where each term of $X$ is finitely generated projective as left $A$-module and as right $B$-module, such that $X\otimes_{B}X^{*}$ and $A$ are isomorphic in $K^{b}(\mbox{$A$-Gr-$A$})$ and $X^{*}\otimes_{A}X$ and $B$ are isomorphic in $K^{b}(\mbox{$B$-Gr-$B$})$. Then $A$ and $B$ are graded Rickard equivalent.
Using [22, Lemmas 3.6 and 3.19], it is straightforward to check the following:
@@
Lemma 4.3 ([26, Lemma 2.6]).
Let $R$ be a commutative ring, $G$ a finite group, $A$ and $B$ fully $G$-graded $R$-algebras. There are mutually inverse functors induce a equivalence between the categories $\Delta(A,B)$-mod and AA-Gr-$B$. In particular, the categories $K^{b}(\mbox{$\Delta(A,B)$-{\rm mod}})$ and $K^{b}(\mbox{$A$-{\rm Gr}-$B$})$ are equivalent.
An explicit pair of mutually inverse functors which induces the equivalence in Lemma 4.3 has been given in [26, Lemma 2.6]. The following theorem is a homotopy category version of [26, Theorem 4.7], which is, as [26, Theorem 4.7], a straightforward consequence of [26, 4.2, 4.3].
@@
Moreover, if the above conditions hold, then $X$ and $X_{1}$ correspond to each other via the category equivalence in Lemma 4.3.
@@
    There is an object $X_{1}\in K^{b}(\Delta(A,B))$ which, regarded as an object of $K^{b}((A_{1}\otimes_{R}B_{1}^{\rm op}))$, induces a Rickard equivalence between $A_{1}$ and $B_{1}$.
Moreover, if above conditions hold, then $X$ and $X_{1}$ are correspond to each other via the category equivalence in Lemma 4.3.
@@
Since Rickard’s paper [30], it has been customary to use the adjective “splendid” to any of the Rickard/derived or Morita equivalences. In [1, Remark 1.2], Boltje defined a more general version of splendid Rickard equivalences, which is more flexible and preserves all the good properties of the original version of Rickard [30] and also later definitions (see for instance [24, Definition 9.7.5]). We adopt Boltje’s definition. If $\varphi:P\xrightarrow{\cong}Q$ is an isomorphism between subgroups $P\leq G$ and $Q\leq H$, we set
@@
    Assume that $p\nmid|G\times H:G_{1}\times H_{1}|$ and $M$ is a/an (complex of) RGRG-RHRH-bimodule(s). If ${\rm Res}_{G_{1}\times H_{1}}^{G\times H}(M)$ is splendid, then $M$ is splendid.
Proof. It suffices to consider bimodules. If $M$ is a $RG_{1}$-$RH_{1}$-bimodule, then a vertex of ${\rm Ind}_{G_{1}\times H_{1}}^{G\times H}(M)$ is contained in a vertex of $M$, and this proves (i). One easily checks statement (ii) by using the Mackey formula. It remains to prove statement (iii). Since $p\nmid|G\times H:G_{1}\times H_{1}|$, $M$ is relatively $G_{1}\times H_{1}$-projective, hence $M$ is isomorphic to a direct summand of
${\rm Ind}_{G_{1}\times H_{1}}^{G\times H}{\rm Res}_{G_{1}\times H_{1}}^{G\times H}(M);$
@@
With the definition of “splendid” in 4.6, tensor products and duals of splendid bimodules/complexes are splendid:
@@
Let $R\in\{\mathcal{O},k\}$. Let $G$, $H$ and $L$ finite groups. If $M$ is a splendid RGRG-RHRH-bimodule and $N$ is a splendid RHRH-RCRC-bimodule, then $M\otimes_{RH}N$ is a splendid RGRG-RLRL-bimodule, and $M^{*}:={\rm Hom}_{R}(M,R)$ is a splendid RHRH-RGRG-bimodule.
@@
Propositions 4.7 (ii) and 4.8 explain why the definition of “splendid” in 4.6 is more flexible than the versions in [30] and [24, Definition 9.7.5].
@@
4.10.
@@
More generally, as in [11, Definition 1.3], we can define splendid modules, complexes, and Rickard equivalences for symmetric interior $P$-algebras:
@@
is a splendid $(A_{1}\otimes_{R}A_{2},B_{1}\otimes_{R}B_{2})$-bimodule. By [22, Lemma 3.6], we have an isomorphism of $(A_{1}\otimes_{R}A_{2},B_{1}\otimes_{R}B_{2})$-bimodules
@@
5 On spin blocks of $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$
Notation 5.1.
Let $n$ be a positive integer. We use $[n]$ to denote the set $\{1,\cdots,n\}$. For conventional convenience we set $0]:=\emptyset$. We denote by $\mathscr{P}(n)$ the set of partitions of $n$ and by $\mathscr{P}_{0}(n)$ the set of strict partitions of $n$, i.e. partitions of $n$ without repeat parts. Set $\mathscr{P}_{0}(0):=\{{\rm empty~partion}\}$. For $\lambda\in\mathscr{P}_{0}(n)$, ${\rm Ab}_{\lambda}$ will denote its $\bar{p}$-abacus ($p$-bar-abacus) display; see [21, §2.3.b] for more details on this. For any non-empty partition $\lambda$, set $h(\lambda):={\rm max}\{k~|~\lambda_{k}>0\}$ to be the length of $\lambda$. We adopt the convention that the 0-th position on the abacus display is always unoccupied, i.e. we always use $h(\lambda)$ beads in ${\rm Ab}_{\lambda}$. Let $\mathscr{P}_{0}:=\bigsqcup_{n\in\mathbb{N}}\mathscr{P}_{0}(n)$. Following [29], we can associate to every $\lambda\in\mathscr{P}_{0}$ its $\bar{p}$-core ${\rm core}(\lambda)\in\mathscr{P}_{0}$. We can also define a non-negative integer called the $\bar{p}$-weight of $\lambda$: ${\rm wt}(\lambda):=(|\lambda|-|{\rm core}(\lambda)|)/p\in\mathbb{N}$.
Assume in this section that $(K,\mathcal{O},k)$ is large enough for $\widetilde{S}_{n}$; this means that $K$ (and hence $\mathcal{O}$) contains a primitive $(2n!)$-th root of unity. We also assume that 1-1 and $2$ have square roots in $K$ (and hence $\mathcal{O}$); this condition is automatic if $n\geq 4$. Let $R\in\{\mathcal{O},k\}$, and $R_{p}\in\{\mathbb{Z}_{p},\mathbb{F}_{p}\}$. If $n>1$, we consider $R\widetilde{S}_{n}$ as a superalgebra corresponding to $(\widetilde{S}_{n})_{\bar{0}}=\widetilde{A}_{n}$; see 3.10. We set
$e_{z}:=(1-z)/2\in R\widetilde{S}_{n}.$
We will use the adjective “spin” when referring to objects associated to $R\widetilde{S}_{n}e_{z}$ rather than the whole of $R\widetilde{S}_{n}$. For example the spin characters of $\widetilde{S}_{n}$ will refer to the $K$-characters $\chi$ of $\widetilde{S}_{n}$ satisfying $\chi(z)=-\chi(1)$, while spin blocks refers to the blocks of $R\widetilde{S}_{n}e_{z}$. Note that $R\widetilde{S}_{n}e_{z}$ inherits the structure of superalgebra from $R\widetilde{S}_{n}$ since the idempotent $e_{z}$ is in $(R\widetilde{S}_{n})_{\bar{0}}$.
Let $G\leq\widetilde{S}_{n}$ with $z\in G$. Following the notation from 3.10, we set

### v5

@@ 5 On Rickard equivalences and splendid complexes
5.1.
Let $R$ be a commutative ring and $G$ a finite group. A finitely generated $R$-algebra AA is called $G$-graded (resp. fully $G$-graded) if, as an $R$-module, AA is a direct sum $A=\oplus_{g\in G}A_{g}$ satisfying $A_{g}A_{h}\subseteq A_{gh}$ (resp. $A_{g}A_{h}=A_{gh}$) for all $g,h\in G$; if $A^{\times}\cap A_{g}\neq\varnothing$, then $A$ is called a crossed product (of $A_{1}$ and $G$). Let $A$, $B$ and $C$ be fully $G$-graded $R$-algebras. A finitely generated AA-$B$-bimodule $M$ is called $G$-graded if, as an $R$-module, $M$ is a direct sum $M=\oplus_{x\in G}M_{x}$ satisfying $A_{g}M_{x}B_{h}=M_{gxh}$ for all $g,h,x\in G$. Hence with this terminology, if $G=\mathbb{Z}/2\mathbb{Z}$, then $A$ and $B$ are exactly $R$-superalgebras and $M$ is exactly an $(A,B)$-bisupermodule. Following [26], we denote by AA-Gr-$B$ the categroy of $G$-graded finitely generated AA-$B$-bimodules and grade preserving AA-$B$-bimodule homomorphisms. Let $K^{b}(\mbox{$A$-Gr-$B$})$ denote the homotopy category of bound complexes over AA-Gr-$B$; see e.g. [24, page 117] for homotopy categories. Let $M$ be a $G$-graded AA-$B$-bimodule and $N$ a $G$-graded $B$-$C$-bimodule. Then $M\otimes_{B}N$ is a $G$-graded AA-$C$-bimodule with $(M\otimes_{B}N)_{x}=\sum_{gh=x}M_{g}\otimes_{B}N_{h}$. Hence for an object $X$ of $K^{b}(\mbox{$A$-Gr-$B$})$ and an object $Y$ of $K^{b}(\mbox{$B$-Gr-$C$})$, the complex $X\otimes_{B}Y$ is an object of $K^{b}(\mbox{$A$-Gr-$C$})$.
Let $A$ and $B$ be symmetric $R$-algebras. If $X$ is a bounded complex of AA-$B$-bimodules that are finitely generated projective as left $A$-modules and as right $B$-modules, such that $X\otimes_{B}X^{*}$ and $A$ are homotopy equivalent as complexes of AA-AA-bimomdules, and that $X^{*}\otimes_{A}X$ and $B$ are homotopy equivalent as complexes of $B$-$B$-bimodules. Then we say that $A$ and $B$ are Rickard equivalent, and $X$ is called a Rickard complex of AA-$B$-bimodules. Here $A$ (resp. $B$) denotes the complex with only one nonzero term $A$ (resp. $B$) in degree 0, and we adopt [24, Definition 1.17.9] for the definition of tensor product of complexes.
Let $A$, $B$ and $C$ be fully $G$-graded symmetric $R$-algebras. If there is a complex (resp. module) $X$ in $K^{b}(\mbox{$A$-Gr-$B$})$ (resp. AA-Gr-$B$) where each term of $X$ is finitely generated projective as left $A$-module and as right $B$-module, such that $X\otimes_{B}X^{*}$ and $A$ are isomorphic in $K^{b}(\mbox{$A$-Gr-$A$})$ (resp. AA-Gr-AA) and $X^{*}\otimes_{A}X$ and $B$ are isomorphic in $K^{b}(\mbox{$B$-Gr-$B$})$ (resp. (resp. $B$-Gr-$B$)). Then $A$ and $B$ are graded Rickard (resp. Morita) equivalent. Note that if $G=\mathbb{Z}/2\mathbb{Z}$, then a graded Morita equivalence is exactly a Morita superequivalence.
Using [23, Lemmas 3.6 and 3.19], it is straightforward to check the following:
@@
The following theorem is a homotopy category version of [26, Theorem 4.7], which is, as [26, Theorem 4.7], a straightforward consequence of [26, 4.2, 4.3].
@@
Moreover, if the above conditions hold, then $X_{1}$ is the $1$-component of $X$, and
$X\cong{\rm Ind}_{\Delta(A,B)}^{A\otimes_{R}B^{\rm op}}(X_{1})\cong A\otimes_{A_{1}}X_{1}\cong X_{1}\otimes_{B_{1}}B.$
@@
    There is an object $X_{1}\in K^{b}(\Delta(A,B))$ which, regarded as an object of $K^{b}(A_{1}\otimes_{R}B_{1}^{\rm op})$, induces a Rickard equivalence between $A_{1}$ and $B_{1}$.
Moreover, if above conditions hold, then $X_{1}$ is the $1$-component of $X$, and
$X\cong{\rm Ind}_{\Delta(A,B)}^{A\otimes_{R}B^{\rm op}}(X_{1})\cong A\otimes_{A_{1}}X_{1}\cong X_{1}\otimes_{B_{1}}B.$
@@
Since Rickard’s paper [30], it has been customary to use the adjective “splendid” to any of the Rickard/derived or Morita equivalences. In [1, Remark 1.2], Boltje defined a more general version of splendid Rickard equivalences, which is more flexible and preserves all the good properties of the original version of Rickard [30] and also later definitions (see for instance [25, Definition 9.7.5]). We adopt Boltje’s definition. If $\varphi:P\xrightarrow{\cong}Q$ is an isomorphism between subgroups $P\leq G$ and $Q\leq H$, we set
@@
    Let $Y$ be a subgroup of $G\times H$ containing $G_{1}\times H_{1}$, and $M$ a/an (complex of) $RY$-module(s). Assume that $p\nmid|Y:G_{1}\times H_{1}|$ and ${\rm Res}_{G_{1}\times H_{1}}^{Y}(M)$ is splendid. Then ${\rm Ind}_{Y}^{G\times H}(M)$ is splendid.
Proof. It suffices to consider bimodules. If $M$ is a $RG_{1}$-$RH_{1}$-bimodule, then a vertex of ${\rm Ind}_{G_{1}\times H_{1}}^{G\times H}(M)$ is contained in a vertex of $M$, and this proves (i). One easily checks statement (ii) by using the Mackey formula. It remains to prove statement (iii). Since $p\nmid|Y:G_{1}\times H_{1}|$, $M$ is relatively $G_{1}\times H_{1}$-projective, hence $M$ is isomorphic to a direct summand of
${\rm Ind}_{G_{1}\times H_{1}}^{Y}{\rm Res}_{G_{1}\times H_{1}}^{Y}(M);$
see e.g. [24, Theorem 2.6.2]. It follows that ${\rm Ind}_{Y}^{G\times H}(M)$ is isomorphic to a direct summand of
${\rm Ind}_{G_{1}\times H_{1}}^{G\times H}{\rm Res}_{G_{1}\times H_{1}}^{Y}(M),$
@@
With the definition of “splendid” in 5.5, tensor products and duals of splendid bimodules/complexes are splendid:
@@
Let $R\in\{\mathcal{O},k\}$. Let $G$, $H$ and $L$ finite groups. If $M$ is a splendid RGRG-RHRH-bimodule and $N$ is a splendid RHRH-RLRL-bimodule, then $M\otimes_{RH}N$ is a splendid RGRG-RLRL-bimodule, and $M^{*}:={\rm Hom}_{R}(M,R)$ is a splendid RHRH-RGRG-bimodule.
@@
Propositions 5.6 (ii) and 5.7 explain why the definition of “splendid” in 5.5 is more flexible than the versions in [30] and [25, Definition 9.7.5].
@@
Theorem 5.9 (a slight generalisation of [30, Theorem 5.2]).
Let $G$ and $H$ be finite groups, $b\in Z(\mathcal{O}G)$ and $c\in Z(\mathcal{O}H)$ be nonzero idempotents. Let $\bar{b}$ and $\bar{c}$ be the image of $b$ and $c$ in $kG$ and $kH$ respectively. If $\bar{X}$ is a splendid Rickard complex of $kG\bar{b}$-$kH\bar{c}$-bimodules. Then there is a splendid Rickard complex $X$ of $\mathcal{O}Gb$-$\mathcal{O}Hc$-bimodules such that $\bar{X}\cong k\otimes_{\mathcal{O}}X$. Moreover, $X$ is unique up to isomorphism.
[…]
The following theorem should be a useful tool for the (refined) Broué conjecture.
Theorem 5.10.
Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$. Assume that $\mathcal{O}^{\prime}$ is finitely generated as $\mathcal{O}$-module. Let $(R^{\prime},R)\in\{(\mathcal{O}^{\prime},\mathcal{O}),(k^{\prime},k)\}$. Let $G$ and $H$ be finite groups. Assume that $K^{\prime}$ contains a primitive $|G\times H|$-th root of unity. Let $b$ be a block of $R^{\prime}G$ and $c$ a block of $R^{\prime}H$. Assume that $b\in RG$ and $c\in RH$. Assume that $RGb$ and $RHc$ are splendidly Rickard equivalent. Then the Brauer correspondent of $RGb$ and the Brauer correspondent of $RHc$ are splendidly Morita equivalent.
Proof. By Theorem 5.9, we may assume that $(R^{\prime},R)=(k^{\prime},k)$. Let $X$ be a splendid Rickard complex of $(kGb,kHc)$-bimodules. Then $X^{\prime}:=k^{\prime}\otimes_{k}X$ is a splendid Rickard complex of $(k^{\prime}Gb,k^{\prime}Hc)$-bimodules. By [2, Theorem 1.6 (a)], the element
$\gamma:=\sum_{i\in\mathbb{Z}}(-1)^{i}[k^{\prime}\otimes_{k}X_{i}]\in T^{\Delta}(k^{\prime}Gb,k^{\prime}Hc)$
is a $p$-permutation equivalence between $k^{\prime}Gb$ and $k^{\prime}Hc$; see [2] for the definition of a $p$-permutation equivalence and the meaning of the notation $T^{\Delta}(k^{\prime}Gb,k^{\prime}Hc)$. Let $(\Delta\varphi,e\otimes f^{*})$ be a maximal $\gamma$-Brauer pair, where $\Delta\varphi$ is a twisted diagonal subgroup of $G\times H$ (with $\varphi:P\xrightarrow{\cong}Q$ an isomorphism between subgroups $P\leq G$ and $Q\leq H$). By [2, Theorem 10.11], $(P,e)$ is a maximal $k^{\prime}Gb$-Brauer pair and $(Q,f)$ is a maximal $k^{\prime}Hc$-Brauer pair. Let $b^{\prime}:=\sum_{g\in[N_{G}(P)/N_{G}(P,e)]}{}^{g}e$ and $c^{\prime}:=\sum_{h\in[N_{H}(Q)/N_{H}(Q,f)]}{}^{h}f$. Then by [25, Theorem 6.7.6 (iii)], $b^{\prime}$ is the Brauer correspondent of $b$ in $k^{\prime}N_{G}(P)$ and $c^{\prime}$ is the Brauer correspondent of $c$ in $k^{\prime}N_{H}(Q)$. Since $b\in kG$ and $c\in kH$, we have $b^{\prime}\in kN_{G}(P)$ and $c^{\prime}\in kN_{H}(Q)$.
Let $M^{\prime}$ be the maximal module of $\gamma$ (see [2, Definition 14.2]). So by [2, Definition 14.2], $M^{\prime}$ is up to isomorphism the unique indecomposable direct summand of the $(k^{\prime}Gb,k^{\prime}Hc)$-bimodule $\bigoplus_{i\in\mathbb{Z}}(k^{\prime}\otimes_{k}X_{i})$ who has vertex $\Delta\varphi$. Then by [20, Lemma 5.1], we deduce that there is an indecomposable direct summand $M$ of the $(kGb,kHc)$-bimodule $\bigoplus_{i\in\mathbb{Z}}X_{i}$ with vertex $\Delta\varphi$, such that $M^{\prime}\cong k^{\prime}\otimes_{k}M$. By [4, (3.4)], the $kN_{G\times H}(\Delta\varphi)$-module ${\rm Br}_{\Delta\varphi}^{k^{\prime}(G\times H)}(M^{\prime})$ is the Green correspondent of the $k^{\prime}(G\times H)$-module $M^{\prime}$. (With the notation of [4], ${\rm Br}_{\Delta\varphi}^{k^{\prime}(G\times H)}(M^{\prime})$ is denoted by $M^{\prime}(\Delta\varphi)$.) Then by [2, Theorem 1.3 (d)], the $p$-permutation $k^{\prime}N_{G}(P)b^{\prime}$-$k^{\prime}N_{H}(Q)c^{\prime}$-bimodule
${\rm Ind}_{N_{G\times H}(\Delta\varphi)}^{N_{G}(P)\times N_{H}(Q)}({\rm Br}_{\Delta\varphi}^{k^{\prime}(G\times H)}(M^{\prime}))$
induces a Morita equivalence between $k^{\prime}N_{G}(P)b^{\prime}$ and $k^{\prime}N_{H}(Q)c^{\prime}$. Now by [20, Proposition 4.5 (c)], the $p$-permutation $kN_{G}(P)b^{\prime}$-$kN_{H}(Q)c^{\prime}$-bimodule
${\rm Ind}_{N_{G\times H}(\Delta\varphi)}^{N_{G}(P)\times N_{H}(Q)}({\rm Br}_{\Delta\varphi}^{k(G\times H)}(M))$
[…]
Remark 5.11.
In fact, Theorem 5.10 holds in greater generality: we can assume that $RGb$ and $RHc$ are $p$-permutation equivalent to obtain the same result. The proof is entirely analogous.
5.12.
@@
More generally, as in [12, Definition 1.3], we can define splendid modules, complexes, and Rickard equivalences for symmetric interior $P$-algebras:
@@
is a splendid $(A_{1}\otimes_{R}A_{2},B_{1}\otimes_{R}B_{2})$-bimodule. By [23, Lemma 3.6], we have an isomorphism of $(A_{1}\otimes_{R}A_{2},B_{1}\otimes_{R}B_{2})$-bimodules
@@
6 On spin blocks of $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$
Notation 6.1.
Let $n$ be a positive integer. We use $[n]$ to denote the set $\{1,\cdots,n\}$. For conventional convenience we set $0]:=\varnothing$. We denote by $\mathscr{P}_{0}(n)$ the set of strict partitions of $n$, i.e. partitions of $n$ without repeat parts. Set $\mathscr{P}_{0}(0):=\{{\rm empty~partion}\}$. For $\lambda\in\mathscr{P}_{0}(n)$, ${\rm Ab}_{\lambda}$ will denote its $\bar{p}$-abacus ($p$-bar-abacus) display; see [22, §2.3.b] for more details on this. For any non-empty partition $\lambda$, set $h(\lambda):={\rm max}\{k~|~\lambda_{k}>0\}$ to be the length of $\lambda$. We adopt the convention that the 0-th position on the abacus display is always unoccupied, i.e. we always use $h(\lambda)$ beads in ${\rm Ab}_{\lambda}$. Let $\mathscr{P}_{0}:=\bigsqcup_{n\in\mathbb{N}}\mathscr{P}_{0}(n)$. Following [29], we can associate to every $\lambda\in\mathscr{P}_{0}$ its $\bar{p}$-core ${\rm core}(\lambda)\in\mathscr{P}_{0}$. We can also define a non-negative integer called the $\bar{p}$-weight of $\lambda$: ${\rm wt}(\lambda):=(|\lambda|-|{\rm core}(\lambda)|)/p\in\mathbb{N}$.
Assume in this section that $(K,\mathcal{O},k)$ is large enough for $\widetilde{S}^{\eta}_{n}$; this means that $K$ (and hence $\mathcal{O}$) contains a primitive $(2n!)$-th root of unity. Following [22, 23] we also assume that 1-1 and $2$ have square roots in $K$ (and hence $\mathcal{O}$); this condition is automatic if $n\geq 4$. Let $(R,R_{p})\in\{(\mathcal{O},\mathbb{Z}_{p}),(k,\mathbb{F}_{p})\}$. If $n>1$, we consider $R\widetilde{S}^{\eta}_{n}$ as a superalgebra corresponding to $(\widetilde{S}^{\eta}_{n})_{\bar{0}}=\widetilde{A}_{n}$; see 4.10. We set
$e_{z}:=(1-z)/2\in R\widetilde{S}^{\eta}_{n}.$
We will use the adjective “spin” when referring to objects associated to $R\widetilde{S}^{\eta}_{n}e_{z}$ rather than the whole of $R\widetilde{S}^{\eta}_{n}$. For example the spin characters of $\widetilde{S}^{\eta}_{n}$ will refer to the $K$-characters $\chi$ of $\widetilde{S}^{\eta}_{n}$ satisfying $\chi(z)=-\chi(1)$, while spin blocks refers to the blocks of $R\widetilde{S}^{\eta}_{n}e_{z}$. Note that $R\widetilde{S}^{\eta}_{n}e_{z}$ inherits the structure of superalgebra from $R\widetilde{S}^{\eta}_{n}$ since the idempotent $e_{z}$ is in $(R\widetilde{S}^{\eta}_{n})_{\bar{0}}$.
Let $G\leq\widetilde{S}^{\eta}_{n}$ with $z\in G$. Following the notation from 4.10, we set